# Compatible Poisson Structures On Multiplicative Quiver Varieties

### v2

@@ 5.3. Integrability
[…]
@@
We can compute in general that $\psi_{\underline{z}}(h_{k;\alpha},h_{l;\beta})=0$ for any indices because $\psi_{\underline{z}}(Z_{ij},-)=0$ and $\psi_{\underline{z}}((W_{\gamma}V_{\gamma})_{ij},-)=0$ by (§ 5.1.1). Thus $\mathcal{H}_{\mathrm{int}}$ and its subalgebra $\mathcal{H}$ stay abelian for any Poisson structure. Using (§ 5.1.1) again, we find as in (5.18),
@@ Appendix A Explicit correspondence
$\langle(\partial_{b})_{ij},\mathrm{d}(\operatorname{Id}+\mathtt{X}_{b}\mathtt{X}_{b^{\ast}})^{\epsilon(b)}_{uv}\rangle$ =$=\left\{\begin{array}[]{lc}\delta_{uj}(\mathtt{X}_{b^{\ast}})_{iv}&\epsilon(b)=+1,\\ -(\operatorname{Id}+\mathtt{X}_{b}\mathtt{X}_{b^{\ast}})^{-1}_{uj}(\mathtt{X}_{b^{\ast}}(\operatorname{Id}+\mathtt{X}_{b}\mathtt{X}_{b^{\ast}})^{-1})_{iv}&\epsilon(b)=-1,\end{array}\right.$ (A.5c)
$\langle(\partial_{b})_{ij},\mathrm{d}(\operatorname{Id}+\mathtt{X}_{b^{\ast}}\mathtt{X}_{b})^{-\epsilon(b)}_{uv}\rangle$ =$=\left\{\begin{array}[]{ll}-((\operatorname{Id}+\mathtt{X}_{b^{\ast}}\mathtt{X}_{b})^{-1}\mathtt{X}_{b^{\ast}})_{uj}(\operatorname{Id}+\mathtt{X}_{b^{\ast}}\mathtt{X}_{b})^{-1}_{iv}&\epsilon(b)=+1,\\ (\mathtt{X}_{b^{\ast}})_{uj}\delta_{iv}&\epsilon(b)=-1,\end{array}\right.$ (A.5f)
@@
(A.4)$_{RHS}$ $=(\partial_{a})_{ij}-\frac{1}{4}\sum_{b=a,a^{\ast}}\Big{[}\mathtt{X}_{a^{\ast}}(\partial_{b}\mathtt{X}_{b}-\mathtt{X}_{b}\partial_{b})(\operatorname{Id}+\mathtt{X}_{a}\mathtt{X}_{a^{\ast}})^{-1}+(\partial_{b}\mathtt{X}_{b}-\mathtt{X}_{b}\partial_{b})(\operatorname{Id}+\mathtt{X}_{a^{\ast}}\mathtt{X}_{a})^{-1}\mathtt{X}_{a^{\ast}}$
@@
$-\frac{1}{2}(\partial_{a})_{ij}+\frac{1}{4}\left[(\operatorname{Id}+\mathtt{X}_{a^{\ast}}\mathtt{X}_{a})^{-1}\partial_{a}(\operatorname{Id}+\mathtt{X}_{a}\mathtt{X}_{a^{\ast}})+(\operatorname{Id}+\mathtt{X}_{a^{\ast}}\mathtt{X}_{a})\partial_{a}(\operatorname{Id}+\mathtt{X}_{a}\mathtt{X}_{a^{\ast}})^{-1}\right]_{ij}\,.$

### v3

@@ 5.3. Integrability
$\mathcal{H}:=$ $\mathbb{C}[h_{k}\mid k\geq 1]\,,\quad\mathcal{Q}:=\mathbb{C}[t_{k;\alpha\beta}\mid k\geq 1,\,\,1\leq\alpha,\beta\leq d]\,$,
$\mathcal{H}_{\mathrm{int}}:=$ $\mathbb{C}[h_{k;\alpha}\mid k\geq 1,\,\,1\leq\alpha\leq d]\,.$
@@
We can compute in general that $\psi_{\underline{z}}(h_{k;\alpha},h_{l;\beta})=0$ for any indices because $\psi_{\underline{z}}(Z_{ij},-)=0$ and $\psi_{\underline{z}}((W_{\gamma}V_{\gamma})_{ij},-)=0$ by (5.9). Thus $\mathcal{H}_{\mathrm{int}}$ and its subalgebra $\mathcal{H}$ stay abelian for any Poisson structure. Using (5.9) again, we find as in (5.18),
@@ Appendix A Explicit correspondence
$\langle(\partial_{b})_{ij},\mathrm{d}(\operatorname{Id}+\mathtt{X}_{b}\mathtt{X}_{b^{\ast}})^{\epsilon(b)}_{uv}\rangle$ $=\left\{\begin{array}[]{lc}\delta_{uj}(\mathtt{X}_{b^{\ast}})_{iv}&\epsilon(b)=+1,\\ -(\operatorname{Id}+\mathtt{X}_{b}\mathtt{X}_{b^{\ast}})^{-1}_{uj}(\mathtt{X}_{b^{\ast}}(\operatorname{Id}+\mathtt{X}_{b}\mathtt{X}_{b^{\ast}})^{-1})_{iv}&\epsilon(b)=-1,\end{array}\right$. (A.5c)
$\langle(\partial_{b})_{ij},\mathrm{d}(\operatorname{Id}+\mathtt{X}_{b^{\ast}}\mathtt{X}_{b})^{-\epsilon(b)}_{uv}\rangle$ $=\left\{\begin{array}[]{ll}-((\operatorname{Id}+\mathtt{X}_{b^{\ast}}\mathtt{X}_{b})^{-1}\mathtt{X}_{b^{\ast}})_{uj}(\operatorname{Id}+\mathtt{X}_{b^{\ast}}\mathtt{X}_{b})^{-1}_{iv}&\epsilon(b)=+1,\\ (\mathtt{X}_{b^{\ast}})_{uj}\delta_{iv}&\epsilon(b)=-1,\end{array}\right$. (A.5f)
@@
[…]
@@
$\frac{1}{2}(\partial_{a})_{ij}+\frac{1}{4}\left[(\operatorname{Id}+\mathtt{X}_{a^{\ast}}\mathtt{X}_{a})^{-1}\partial_{a}(\operatorname{Id}+\mathtt{X}_{a}\mathtt{X}_{a^{\ast}})+(\operatorname{Id}+\mathtt{X}_{a^{\ast}}\mathtt{X}_{a})\partial_{a}(\operatorname{Id}+\mathtt{X}_{a}\mathtt{X}_{a^{\ast}})^{-1}\right]_{ij}\,.$